# Challenging Excited States from Adaptive Quantum Eigensolvers: Subspace Expansions vs. State-Averaged Strategies

### v1

@@ 4. Conclusions
Although q-sc-EOM fails to recover correct energy curves with small excitation manifolds, even with a converged ground state, it is unclear to what extent MORE-ADAPT-VQE will require longer circuits. To obtain $k$ different states, MORE-ADAPT-VQE will naïvely need to perform $k$ times more measurements than traditional ADAPT-VQE at each iteration. For intensive properties like excitation energies, $k$ can be constant with respect to system size. We have seen in section 3 that MORE-ADAPT-VQE does not seem to require more than $k$ references to obtain these $k$ excited states. Our results suggest that this is not the case for q-sc-EOM, however. If the full set of singles and doubles used in the original paper[26] is needed as an expansion basis, the construction of $\bar{H}$ will require $\mathcal{O}\left(N^{8}\right)$ measurements as a function of system size $N$. Testing on larger systems is needed to meaningfully compare the necessary sizes of the expansion manifolds for each method, but the current data suggests that MORE-ADAPT-VQE achieves better accuracy with the same number of references. Like q-sc-EOM, MORE-ADAPT-VQE has the benefit of not computing an overlap matrix, which has been shown to substantially improve resilience to noise. Furthermore, it requires no additional qubits compared to ground state methods, provided linear combinations of reference pairs are easily prepared, utilizing the off-diagonal measurement scheme from Ref. 26.
@@
for any number of orthogonal references {|$\{\ket{\phi_{i}}\}$ and orthogonal target states {|$\{\ket{\Psi_{i}}\}$? If this condition is satisfied, it follows that MORE-ADAPT-VQE can, in the absence of ansatz minima, obtain the exact eigenstates of the Hamiltonian. A stronger requirement is that:
@@
for all $i$, with the result that properties of individual states would no longer require off-diagonal measurements of operators, such as the electric dipole moment. Proving or disproving these conjectures would offer significant insight into the potential and limitations of MORE-ADAPT-VQE.
@@ 9. Author contributions
Both authors contributed to the development of the theory. Multistate contracted VQE, MORE-ADAPT-VQE, and q-sc-EOM codes were implemented by H.R.G. Data were obtained by H.R.G. Both authors contributed to manuscript preparation.

### v2

@@ 4. Conclusions
To obtain $k$ different states, MORE-ADAPT-VQE will naïvely need to perform $k$ times more measurements than traditional ADAPT-VQE at each iteration, since the $k^{2}$ elements of $\bar{H}$ need only be measured at the very end of the algorithm. This assumes that the size of $k$ is smaller than the size of the pool, which is a realistic expectation. In the operator addition step of ADAPT-VQE, a gradient element must be computed for every element in the pool. In the UCCGSD pool, for example, this corresponds to $\mathcal{O}\left(N^{4}\right)$ gradients, with an additional factor of $k$ from the different references in MORE-ADAPT-VQE, while $\bar{H}$ only has $k^{2}$ elements to measure. For intensive properties like excitation energies, $k$ can be constant with respect to system size. We have seen in section 3 that MORE-ADAPT-VQE does not seem to require more than $k$ references to obtain these $k$ excited states. This would imply that the dimension of $\bar{H}$ does not need to increase just because the Hilbert space becomes combinatorially larger with respect to system size. Our results suggest that this is not the case for q-sc-EOM, however. If the full set of singles and doubles used in the original paper[26] is needed as an expansion basis, the construction of $\bar{H}$ will require $\mathcal{O}\left(N^{8}\right)$ measurements as a function of system size $N$. Testing on larger systems is needed to meaningfully compare the necessary sizes of the expansion manifolds for each method, but the current data suggests that MORE-ADAPT-VQE achieves better accuracy with the same number of references. Like q-sc-EOM, MORE-ADAPT-VQE has the benefit of not computing an overlap matrix, which has been shown to substantially improve resilience to noise. Furthermore, it requires no additional qubits compared to ground state methods, provided linear combinations of reference pairs are easily prepared, utilizing the off-diagonal measurement scheme from Ref. 26. Although q-sc-EOM fails to recover correct energy curves with small excitation manifolds, even with a converged ground state, it is unclear to what extent MORE-ADAPT-VQE will require longer circuits, which could prove to be a limiting factor.
@@
for any number of orthogonal references {|$\{\ket{\phi_{i}}\}$ and orthogonal target states {|$\{\ket{\Psi_{i}}\}$? If this condition is satisfied, it follows that MORE-ADAPT-VQE can, in the absence of ansatz minima, obtain the exact eigenstates of the Hamiltonian. A stronger requirement is that
@@
for all $i$, with the result that properties of individual states would no longer require off-diagonal measurements of operators, such as the electric dipole moment. Addressing these conjectures would offer significant insight into the potential and limitations of MORE-ADAPT-VQE.
In this work, we have exclusively used the naïve choice of equal weights. References [66] and [67] examine state-averaged cost functions with different weight schemes, as in SS-VQE, providing theoretical justification for using specific sets of weights depending on what cost function one is trying to minimize. For example, different sets of weights are expected to give better estimates of the wavefunctions than the weights that give the best estimates of the energies. Using alternative weights merits exploration in the context of MORE-ADAPT-VQE. As reference [67] points out, the trace of $\bar{H}$ is invariant to rotations among the references {|$\left\{\ket{\phi_{i}}\right\}$, meaning that for equal weights, minimizing the average expectation value of $\bar{H}$ in the basis of the references,
$\sum_{i}\omega_{i}E_{i},$ (19)
is equivalent to minimizing the same average expectation value in the basis of the Ritz eigenstates,
$\sum_{i}\omega_{i}E_{i}^{\prime}.$ (20)
This is not true for the case of unequal weights. This discrepancy highlights a choice we have made in defining MORE-ADAPT-VQE to minimize the former quantity rather than the latter. In minimizing the former quantity, we would be pursuing the “wrong” objective function to some extent, though this is much cheaper than computing roughly $k^{2}$ off-diagonal gradient elements for each operator in each iteration, necessary for the latter possibility. We explore the problems of weight selection and cost function definition in a manuscript currently in preparation.
@@ 9. Author contributions
Both authors contributed to the development of the theory. MC-VQE, MORE-ADAPT-VQE, and q-sc-EOM codes were implemented by H.R.G. Data were obtained by H.R.G. Both authors contributed to manuscript preparation.